# The Quantum Decoding Problem : Tight Achievability Bounds and Application to Regev’s Reduction

### v1

@@ 1.1 State of the art
erase the 11st register by decoding $\boldsymbol{c}$}: $\mapsto$ $\quad|{\boldsymbol{0}}\rangle\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}+\boldsymbol{e}\rangle$
@@
It is worthwhile to note that by using the same decoding algorithm with a slight twist in Regev’s reduction, we can construct $\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ for any $\boldsymbol{v}\in\mathbb{F}_{q}^{n}$. Indeed, Regev’s reduction is slightly modified as follows
$\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$
@@
$\xrightarrow{\textrm{Decoder}}$ $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}+\boldsymbol{e}\rangle$
@@
$\xrightarrow{\operatorname{QFT}}$ $\frac{q^{n}}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle.$
@@
By measuring the resulting state $\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ we obtain an element in $\boldsymbol{v}+{\mathscr{C}}^{\perp}$ such that $\boldsymbol{y}$} is relatively small. In other words, we solve the variant of the decoding problem for the dual code $\mathscr{C}$$^{\perp}$ for which we just have to find for any word $\boldsymbol{v}$ of the ambient space a sufficiently close codeword. This is generally the decoding problem as it appears in rate distortion theory [CT91, Chap 10]. It might be also called “decoding in the surjective regime”, where we are not asked to find the closest codeword, but a close enough codeword. In other words, Regev’s reduction can be tweaked to transform a classical decoding algorithm working in the injective regime into a quantum decoding algorithm working in the surjective regime. Recently this approach has been applied on structured code families such as Reed-Solomon codes [JSW+24, CT25] for which there exist efficient decoding algorithms. This approach yields a polynomial time algorithm for decoding the dual Reed Solomon code in the surjective regime. For suitable choices of the function $f$ this seems to yield a quantum polynomial time algorithm for decoding in cases where the best known classical algorithms take exponential time, thus yielding a problem which is a good candidate for quantum advantage.
@@
Problem 3 (Quantum Decoding Problem $\text{QDP}(q,n,k,f)$).
@@
Note that this problem reduces to the classical decoding problem $\mathrm{DP}(q,n,k,p)$ where $p({\boldsymbol{x}})=|f({\boldsymbol{x}})|^{2}$ by measuring the quantum input state. However, this quantum decoding problem might turn out to be intrinsically easier than classical decoding. Indeed, for an $f$ which has a product structure $f(\boldsymbol{e})=\prod_{i=1}^{n}g(e_{i})$, Chen, Liu and Zhandry showed in [CLZ22] that one can use a well chosen quantum measurement on each qudit of $|\psi_{\boldsymbol{c}}\rangle$ (note that in this case $|\psi_{\boldsymbol{c}}\rangle$ has the form $\bigotimes_{i=1}^{n}\sum_{e_{i}\in{\mathbb{F}_{q}}}g(e_{i})|c_{i}+e_{i}\rangle$) yielding partial information on the symbols $c_{i}$ of the codeword $\boldsymbol{c}$}. This in turn can be used together with the Arora-Ge algorithm [AG11] to recover $\boldsymbol{c}$}. With a well chosen function $g$ which is such that its Fourier transform $\hat{g}$ has support included in $\{-\frac{q-1}{2}+c,\cdots,\frac{q-1}{2}-c\}$ for some constant $c$, one obtains after performing the rest of Regev’s reduction (namely applying the QFT after erasing the first register and measuring at the end) codewords $\boldsymbol{y}$} in $\mathscr{C}$$^{\perp}$ which are of infinite norm $\lvert\kern-1.0pt\lvert{\boldsymbol{y}}\rvert\kern-1.0pt\rvert_{\infty}\leq\frac{q-1}{2}-c$. It is then shown in [CLZ22] that this results for constant $c$ in a quantum polynomial time algorithm performing this task of producing rather low weight codewords for the infinite norm in a regime of parameters for which all known classical algorithms have subexponential complexity. In other words, this gives an excellent candidate for providing a quantum advantage. This research thread has been further explored in [CHL+23, DFS24].
@@ 1.2 Contributions
Problem 4 (Quantum Decoding Problem $\text{QDP}(\mathscr{C},f)$).
@@
The strong converse theorem [ON00] already shows that ${p_{\textrm{succ}}}=\operatorname{negl}(n)$, however the proof technique of Theorem 2 is of independent interest. The proof for the case at hand is much shorter. We will make use of the particular form of the quantum channel states $\rho_{\alpha}$ at hand. In terms of proof techniques, we look again at the quantum Fourier transforms of our states $\widehat{|\psi_{\boldsymbol{c}}\rangle}$ and, by using the notion of typical sets, we can show that all of the states $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ approximately lie in the same subspace of dimension $q^{nH_{q}(|\widehat{g}|^{2})}$. The condition $H_{q}(|\widehat{g}|^{2})<R$ implies $\frac{q^{nH_{q}(|\widehat{g}|^{2})}}{q^{Rn}}=\operatorname{negl}(n)$. Since there are $q^{Rn}$ different codewords, we will be able to recover any $\boldsymbol{c}$} from $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ only with negligible probability. Formally, we perform this argument using quantum information theory techniques which are inspired by conditional quantum min-entropy notions (see [KRS09] for example).
@@
In all our proofs we make heavily use of the notion of typical set associated to the probability distribution $p\mathop{=}\limits^{\triangle}|\widehat{f}|^{2}$. This notion applies to more general probability distributions than those corresponding to memoryless c-q channels. We have given a more general form, namely Theorems 4 and 5 giving a tight achievability in the case where for the probability distribution $p$ the elements of probability $\approx q^{H_{q}(p)}$ carry almost all the probability of the space (meaning that they form a typical set in the usual sense). In this case again, the critical rate marking the distinction between achievability of decoding and non achievability corresponds to the entropy per symbol $\frac{H_{q}(p)}{n}$. We illustrate this behavior in a case which is of interest in code based cryptography which corresponds to the rank metric. This alternative in code-based cryptography to the usual Hamming metric has lead to submissions to the NIST competition for post-quantum cryptography [ABD+19, AAB+24] which compare favorably to the corresponding submissions in Hamming metric. It can be used to obtain for instance efficient public key encryption or key exchange schemes [ABD+19, ADG+24], signature schemes [AAB+24, BCF+25], or somewhat homomorphic encryption schemes [ADG25]. In this case, the distribution of $p$ does not correspond to a memoryless channel. Nevertheless, we give a tight achievability result in this case (see Theorems 6 and 7) as a corollary of Theorems 4 and 5. The probability distribution $p$ used in this case is a decreasing function of the rank weight. We also give a more general form of Theorem 3 showing that Regev’s reduction when applied to a variant of the PGM yields elements of the dual code of maximal probability at the limit where the PGM works. This more general result applies also to the rank metric case and shows that again we get at the limit where the PGM works elements in the dual code of minimum rank weight.
@@ 2.2 Information Theory
We will use this notion to define what is a typical set for a probability distribution defined over $\mathbb{F}_{q}^{n}$:
@@ 3.1 Notation and Preliminaries
The Holevo capacity $C_{{\mathcal{H}}}$ of this pure state channel is equal to $\sup_{p_{\alpha}}S\left(\sum_{\alpha\in\mathbb{F}_{q}}p_{\alpha}\rho_{\alpha}\right)$ where
@@
$(1-\delta)\log_{q}(B_{i}(\epsilon))\leq H_{q}(p_{i})\leq-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta n_{i},$ (7)
@@
$\leq$ $-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta\left(\sum_{{\boldsymbol{x}}\in\mathbb{F}_{q}^{n_{i}}\setminus{{\mathcal{T}}_{\epsilon}^{(n_{i})}}}-\log_{q}(q_{i}({\boldsymbol{x}}))q_{i}({\boldsymbol{x}})\right)+\delta\log_{q}(\delta)$
$\leq$ $-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta n_{i}.$
The last inequality comes from the fact that $q_{i}$ can be interpreted as a probability distribution over $\mathbb{F}_{q}^{n_{i}}\setminus{{\mathcal{T}}_{\epsilon}^{(n_{i})}}$ and therefore $\log_{q}(q_{i}({\boldsymbol{x}}))q_{i}({\boldsymbol{x}})=H_{q}(q_{i})$ and by using that the $q$-ary entropy of a random variable taking $M$ different values is upper-bounded by $\log_{q}(M)$.
In the case of a nice typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} (7) gives

### v2

@@ 1.1 State of the art
[…]
@@
It is worthwhile to note that by using the same decoding algorithm with a slight twist in Regev’s reduction, we can construct $\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ for any $\boldsymbol{v}\in\mathbb{F}_{q}^{n}$. Indeed, Regev’s reduction is slightly modified by starting instead of the state $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$ by a slight twist on the state obtained by tweaking the amplitudes with characters as follows $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$. From this state, we perform now the same set of operations as in Regev’s reduction:
$\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$ (1)
@@
$\xrightarrow{\textrm{Decoder}}$ $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}+\boldsymbol{e}\rangle$ (2)
@@
$\xrightarrow{\operatorname{QFT}}$ $\frac{q^{n}}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle.$ (3)
By measuring the resulting state $\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ we obtain an element in $\boldsymbol{v}+{\mathscr{C}}^{\perp}$ such that $\boldsymbol{y}$} is relatively small. In other words, we solve the variant of the decoding problem for the dual code $\mathscr{C}$$^{\perp}$ for which we just have to find for any word $\boldsymbol{v}$ of the ambient space a sufficiently close codeword. This is given here by
Problem 3 (Close Codeword Finding in the surjective regime (or finding a close enough codeword) $\mathrm{DP}(q,n,k,w)$).
$q,n,k,w$ are positive integers such that $k<n$. $\mathbf{H}$ is chosen uniformly at random in $\mathbb{F}_{q}^{(n-k)\times n}$ and $\boldsymbol{y}$} uniformly at random in $\mathbb{F}_{q}^{n}$
Given: $\mathbf{H}$, $\boldsymbol{y}$},
[…]
@@
It generally appears like that in rate distortion theory, but in this case the code is generally not chosen uniformly at random but in a family where the problem of finding a close enough codeword can be easily solved usch as for instance convolutional codes, polar codes or LDGM codes. It might be also called “decoding in the surjective regime”, where we are not asked to find the closest codeword, but a close enough codeword. In other words, Regev’s reduction can be tweaked to transform a classical decoding algorithm working in the injective regime into a quantum decoding algorithm working in the surjective regime. Recently this approach has been applied on structured code families such as Reed-Solomon codes [JSW+25, CT25] for which there exist efficient decoding algorithms in the injective regime but not in the surjective regime. This approach yields a polynomial time algorithm for decoding the dual Reed Solomon code (which is also a Reed-Solomon code in that case since it is chosen as having full support) in the surjective regime. For suitable choices of the function $f$ this seems to yield a quantum polynomial time algorithm for decoding in cases where the best known classical algorithms take exponential time, thus yielding a problem which is a good candidate for quantum advantage.
@@
Problem 4 (quantum decoding problem $\text{QDP}(q,n,k,f)$).
@@
Note that this problem reduces to the classical decoding problem $\mathrm{DP}(q,n,k,p)$ where $p({\boldsymbol{x}})=|f({\boldsymbol{x}})|^{2}$ by measuring the quantum input state. However, this quantum decoding problem might turn out to be intrinsically easier than classical decoding. Indeed, for an $f$ which has a product structure $f(\boldsymbol{e})=\prod_{i=1}^{n}g(e_{i})$, Chen, Liu and Zhandry showed in [CLZ22] that one can use a well chosen quantum measurement on each qudit of $|\psi_{\boldsymbol{c}}\rangle$ (note that in this case $|\psi_{\boldsymbol{c}}\rangle$ has the form $\bigotimes_{i=1}^{n}\sum_{e_{i}\in{\mathbb{F}_{q}}}g(e_{i})|c_{i}+e_{i}\rangle$) yielding partial information on the symbols $c_{i}$ of the codeword $\boldsymbol{c}$}. This in turn can be used together with the Arora-Ge algorithm [AG11] to recover $\boldsymbol{c}$}. With a well chosen function $g$ which is such that its Fourier transform $\hat{g}$ has support included in $\{-\frac{q-1}{2}+c,\cdots,\frac{q-1}{2}-c\}$ for some constant $c$, one obtains after performing the rest of Regev’s reduction (namely applying the QFT after erasing the first register and measuring at the end) codewords $\boldsymbol{y}$} in $\mathscr{C}$$^{\perp}$ which are of infinite norm $\lvert\kern-1.0pt\lvert{\boldsymbol{y}}\rvert\kern-1.0pt\rvert_{\infty}\leq\frac{q-1}{2}-c$. It is then shown in [CLZ22] that this results for constant $c$ in a quantum polynomial time algorithm performing this task of producing rather low weight codewords for the infinite norm in a regime of parameters for which all classical algorithms known at that time had subexponential complexity. In other words, this gave an excellent candidate for providing a quantum advantage. This research thread has been further explored in [CHL+23, DFS24]. Note that very recently a new classical algorithm was devised [KOW25] for finding low weight codewords for the infinite norm. It outputs in polynomial a solution of this problem for a slightly larger set of parameters for which the polynomial quantum algorithm of [CLZ22] works, thus showing that this quantum algorithm does not yield a quantum speedup anymore. In any case, this whole thread of research has definitely proved to be fruitful for improving significantly the state of the art for finding short codewords for the infinite norm.
@@ 1.2 Contributions
This theorem applies to a memoryless channel, but it turns out that a more general result is true which applies to more general noise models, namely Theorem 5 and Theorem 1 is only a corollary of this more general result.
@@
Problem 5 (quantum decoding problem $\text{QDP}(\mathscr{C},f)$).
@@
The strong converse theorem [ON00] already shows that ${p_{\textrm{succ}}}=\operatorname{negl}(n)$, however the proof technique of Theorem 2 is of independent interest. The proof for the case at hand is much shorter. We will make use of the particular form of the quantum channel states $\rho_{\alpha}$ at hand. In terms of proof techniques, we look again at the quantum Fourier transforms of our states $\widehat{|\psi_{\boldsymbol{c}}\rangle}$ and, by using the notion of typical sets, we can show that all of the states $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ approximately lie in the same subspace of dimension $q^{nH_{q}(|\widehat{g}|^{2})}$. The condition $H_{q}(|\widehat{g}|^{2})<R$ implies $\frac{q^{nH_{q}(|\widehat{g}|^{2})}}{q^{Rn}}=\operatorname{negl}(n)$. Since there are $q^{Rn}$ different codewords, we will be able to recover any $\boldsymbol{c}$} from $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ only with negligible probability. Formally, we perform this argument using quantum information theory techniques which are inspired by conditional quantum min-entropy notions (see [KRS09] for example). Again, Theorem 2 follows from a more general result, namely Theorem 6, which applies to more general noise models which are non necessarily memoryless and which follows from considerations on the typical set of errors output by the channel.
@@
In all our proofs we make heavily use of the notion of typical set associated to the probability distribution $p\mathop{=}\limits^{\triangle}|\widehat{f}|^{2}$. This notion applies to more general probability distributions than those corresponding to memoryless c-q channels. We have given a more general form, namely Theorems 5 and 6 giving a tight achievability in the case where for the probability distribution $p$ the elements of probability $\approx q^{H_{q}(p)}$ carry almost all the probability of the space (meaning that they form a typical set in the usual sense). In this case again, the critical rate marking the distinction between achievability of decoding and non achievability corresponds to the entropy per symbol $\frac{H_{q}(p)}{n}$. We illustrate this behavior in a case which is of interest in code based cryptography which corresponds to the rank metric. This alternative in code-based cryptography to the usual Hamming metric has lead to submissions to the NIST competition for post-quantum cryptography [ABD+19, AAB+24] which compare favorably to the corresponding submissions in Hamming metric. It can be used to obtain for instance efficient public key encryption or key exchange schemes [ABD+19, ADG+24], signatu re schemes [AAB+24, BCF+25], or somewhat homomorphic encryption schemes [ADG25]. In this case, the distribution of $p$ does not correspond to a memoryless channel. Nevertheless, we give a tight achievability result in this case (see Theorems 7 and 8) as a corollary of Theorems 5 and 6. The probability distribution $p$ used in this case is a decreasing function of the rank weight. We also give a more general form of Theorem 3 showing that Regev’s reduction when applied to a variant of the PGM yields elements of the dual code of maximal probability at the limit where the PGM works. This more general result applies also to the rank metric case and shows that again we get at the limit where the PGM works elements in the dual code of minimum rank weight.
The decoding problem in the surjective regime and an operational version of the duality result of J. Renes [Ren18]
If we replace in Regev’s reduction procedure the initial state $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$ by $\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}},\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle$, solve the quantum decoding problem with the PGM as in Algorithm 4 (for more details see §4.4), then we output elements $\boldsymbol{v}+{\mathscr{C}^{\perp}}$ of maximal probability $p$ at the limit where PGM works. More precisely we obtain
Theorem 4.
Let $g:\mathbb{F}_{q}\rightarrow\mathbb{C}$ with $\lvert\kern-1.0pt\lvert g\rvert\kern-1.0pt\rvert_{2}=1$. Let $R$ be a real number in $(0,H_{q}(|\widehat{g}|^{2}))$. Let $f:\mathbb{F}_{q}^{n}\rightarrow\mathbb{C}$ be so that $f=g^{\otimes n}$. Let $\boldsymbol{v}$ be an element in $\mathbb{F}_{q}^{n}$. Let $\mathscr{C}$} be a linear code on $\mathbb{F}_{q}^{n}$ with generator matrix $\mathbf{G}$ drawn uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn\rfloor\times n}$. For any $\epsilon>0$:
  • Algorithm 4.4 outputs elements ${\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}^{\perp}}$ such that $p({\boldsymbol{y}})\geq q^{-n(H_{q}(|\widehat{g}^{2}|)-\epsilon)}$ with probability $1-\operatorname{negl}(n)$.
  • The probability over $\mathbf{G}$ that there exists an element $\boldsymbol{y}$} in $\boldsymbol{v}+{\mathscr{C}^{\perp}}$ such that $p({\boldsymbol{y}})$ is larger than $q^{-n(R-\epsilon)}$ is a negligible function of $n$.
This gives an optimal lossy compression procedure for the dual code $\mathscr{C}^{\perp}$} based on a optimal decoding procedure based on PGM for $\mathscr{C}$} for the classical-quantum c-q channel corresponding to $g$. In some sense it can be viewed (with the exception of the $\sqrt{2\epsilon}$ error term which appears in [Ren18, §5, Cor. 2]) as an explicit version of the lossy compression procedure mentioned in [Ren18, §5, Cor. 2].
@@ 2.2 Information Theory
We use this notion to define the typical set of a probability distribution over $\mathbb{F}_{q}^{n}$:
@@ 3.1 Notation and Preliminaries
The Holevo capacity $C_{{\mathcal{H}}}$ of this pure state channel is equal to $\sup_{p_{\alpha}}S\left(\sum_{\alpha\in\mathbb{F}_{q}}p_{\alpha}\rho_{\alpha}\right)$ where the supremum is taken over all probability distributions $(p_{\alpha})_{\alpha\in\mathbb{F}_{q}}$ over $\mathbb{F}_{q}$ and
@@
$(1-\delta)\log_{q}(B_{i}(\epsilon))\leq H_{q}(p_{i})\leq-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta n_{i}-\delta\log\delta,$ (10)
@@
$\leq$ $-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta\left(\sum_{{\boldsymbol{x}}\in\mathbb{F}_{q}^{n_{i}}\setminus{{\mathcal{T}}_{\epsilon}^{(n_{i})}}}-\log_{q}(q_{i}({\boldsymbol{x}}))q_{i}({\boldsymbol{x}})\right)-\delta\log_{q}(\delta)$
$\leq$ $-(1-\delta)\log_{q}(A_{i}(\epsilon))+\delta n_{i}-\delta\log_{q}(\delta).$
The last inequality comes from the fact that $q_{i}$ can be interpreted as a probability distribution over $\mathbb{F}_{q}^{n_{i}}\setminus{{\mathcal{T}}_{\epsilon}^{(n_{i})}}$ and therefore $\sum_{{\boldsymbol{x}}\in\mathbb{F}_{q}^{n_{i}}\setminus{{\mathcal{T}}_{\epsilon}^{(n_{i})}}}\log_{q}(q_{i}({\boldsymbol{x}}))q_{i}({\boldsymbol{x}})=H_{q}(q_{i})$ and by using that the $q$-ary entropy of a random variable taking $M$ different values is upper-bounded by $\log_{q}(M)$.
In the case of a nice typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$}, (10) gives